1. A parallel circuit contains four 1.5 volt batteries and three resistances of 5.0, 2.5 and 4.0 - What is the total emf?
  2. What is the total resistance of this circuit?
  3. What current does the parallel circuit above draw?


Thank you


1
Expert's answer
2021-04-13T06:39:12-0400


(1)


"E_{res}=\dfrac{\dfrac{E_1}{R_1}+\dfrac{E_2}{R_2}+\dfrac{E_3}{R_3}}{\dfrac{1}{R_1}+\dfrac{1}{R_2}+\dfrac{1}{R_3}}"

"E_{res}=\dfrac{1.5(\dfrac{1}{R_1}+\dfrac{1}{R_2}+\dfrac{1}{R_3})}{(\dfrac{1}{R_1}+\dfrac{1}{R_2}+\dfrac{1}{R_3})}=1.5\space V"


(2) "\dfrac{1}{R_{res}}=\dfrac{1}{R_1}+\dfrac{1}{R_2}+\dfrac{1}{R_3}"


"\Rightarrow \dfrac{1}{R_{res}}=\dfrac{1}{5}+\dfrac{1}{2.5}+\dfrac{1}{4}"

"\Rightarrow R_{res}=\dfrac{20}{17}=1.17\space\Omega"


(3) "I_{res}=\dfrac{E_{res}}{R_{res}}=\dfrac{1.5}{1.17}"

"\Rightarrow I_{res}=1.282\space A"